# Weak Memory Model Formalisms: Introduction and Survey

### v1

@@ 6.3. Axiomatic Formalisms
One of the most successful approaches to formalising memory models is via constraints on the allowed behaviours. This is referred to as the axiomatic (or declarative) approach in the literature [36; 263]. Given a set of memory events (where each read has a return value), one starts with all possible permutations, and retains only those permutations that satisfy the axioms of the particular memory model (conceptually the opposite of first considering the set of behaviours allowed under Sequential Consistency and then expanding that set to capture the weak behaviours).
Due to its preciseness and relative conciseness, the axiomatic approach is widely adopted for the major memory models, e.g., x86-TSO [227], Arm [27], RISC-V [249], Java [208], and C [50].
Some of the most influential tools in this area are the domain-specific language Cat and its associated litmus-test simulator Herd, introduced in [36]. Cat and Herd are utilised by Arm to formalise its memory model [27], and Cat is also used in a number of model checkers for expressing other memory models (see Section 8.1). An axiomatic formalisation is leveraged by the Check suite of tools and techniques for investigating correctness and security of microarchitectural implementations [197; 204; 283; 205; 199; 284].
The downside of axiomatic formalisms is that they are based on global reasoning about completed executions, where every has a value, and the relations and are supplied (see Section 5). As such, axiomatic formalisms tend not to be suited for reasoning about programs compositionally; that is, one cannot form the set of behaviours of a system by considering individual processes in isolation. The purely axiomatic approach has less applicability to theorem proving (however, see [26; 143]), so operational semantics (see Section 6.2) are sometimes developed in tandem, often with an underlying axiomatic flavour [110; 165; 241; 95].
@@ 6.4. Event Structures and Pomsets
Both event structures [294] and pomsets [237] were invented in the 1980s as mathematical models of general concurrency. An early use of event structures for weak memory models is by Cenciarelli et al. [77] in 2007, who described a small fragment of the Java memory model by combining event structures with operational and axiomatic techniques. Then from 2016 onwards, there have been many attempts to define memory models using related ideas [74; 155; 232; 79; 231; 153; 156; 218]. The Weakestmo Semantics and its extension [79; 218] are multi-execution, i.e., they encode multiple execution graphs in a single event structure (cf. the Promising Semantics Section 6.2).
@@ 6.5. Program Transformations and Instruction Reorderings
A number of early works [17; 210; 211] specified weak memory models in terms of which pairs of instructions can be reordered. This aligns with the workings of the processor pipeline (Definition 2.1), and is often the level at which processors are described informally. The first formal memory-model framework based on instruction reordering was developed by Arvind and Maessen [41], who included store-atomicity as an extra criterion.
Possible instruction reordering is important in the context of compiler-related aspects of weak memory models. Common program transformations (including instruction reordering) performed by optimising compilers are not always correct under weak memory models [258; 288]. To formulate what it means for a program transformation to be “correct” under a memory model, one typically develops some trace-based denotational semantics, and a transformation is then correct if the denotation of the transformed program is (a subset of) the denotation of the original program. Works in this line of research include Ševčík and Aspinall [258] and Burckhardt et al. [69] for the Java memory model, and Dodds et al. [109] and Dvir et al. [117] for the C memory model and its RA fragment.
Instruction reordering was also the basis in [85], where the memory model is encoded as a parameter to sequential composition on the level of the program text. Weak memory effects are explained in terms of the standard operators for (no-parameter) sequential composition and parallel composition, via a set of transformation rules on the instructions, in the spirit of the Concurrent Kleene Algebra [149; 146]. For example, instantiating the SB litmus test on an x86 (TSO) machine, one can derive the following.
@@
This reasoning framework has been used as the basis for addressing security vulnerabilities related to speculative execution [91] and information flow [267]. By transforming a program under a weak memory model to a program under Sequential Consistency, standard reasoning methods and tools become amenable, saving the need for specialised logics and techniques. However, in contrast to timestamped messages, these program-level transformations cannot explain non-multicopy-atomic behaviours [182].
@@ 7. Program Logics and mechanised support
There are three prominent families of program logics for concurrent programs that have been adapted to reason about weak memory effects: Owicki-Gries (OG) [228], Rely-Guarantee (RG) [157; 158], and Concurrent Separation Logic (CSL) [152; 247; 64]. This section surveys these adaptations; a summary is given in Table 5.
@@
| OG/RG | Axiomatic | C fragment | RCU | [181] |
| RG | Operational | x86-TSO | Simpson’s | [248] |
| RG | Operational | (parameterised) | litmus tests | [13] |
| OG | View-based | C fragment | Peterson’s, RCU | [96; 99; 110; 295] |
| RG | potential-based | Causal Consistency | litmus tests, Peterson’s | [179] |
| CSL | Axiomatic | C fragment | MS queue and others | [112; 113; 289; 286] |
| CSL | Operational | C fragment | MS queue and others | [164] |
| CSL | Operational | Rust | Rust library | [102] |
| CSL | Operational | OCaml | locks, bounded queue | [217; 216] |
| RG | Reordering | hardware models | Chase-Lev, Peterson’s | [93; 94] |
| Invariant | Axiomatic | (parameterised) | PostgreSQL | [30; 26] |
| Invariant | Operational | C fragment | Peterson’s | [111] |
@@
Based on an axiomatic formulation of the release-acquire fragment of the C memory model, a rely-guarantee logic was developed and applied to the read-copy-update synchronisation mechanism [181].
@@
A rely-guarantee logic on the x86 assembly instructions was developed in [248], based on the operational formulation of x86-TSO [227]. This logic was applied to Simpson’s four-slot algorithm, and mechanised in HOL. The rely-guarantee logic of [13] works on general memory models that are specified operationally, with case-studies performed on litmus tests.
Another line of research [96; 99] devised a view-based formalisation (see Section 6.2) of the release-acquire-relaxed fragment of the C memory model, and built an Owicki-Gries logic based on this formulation. The main case-studies, carried out in Isabelle/HOL, are Peterson’s algorithm and the read-copy-update mechanism; related works in this line of research include [110; 295].
Similar to view-based semantics is the “potential-based” semantics, which has been used as the basis of a rely-guarantee logic for Causal Consistency [179].
@@
One line of research consists of the following three generations of logics: RSL (Relaxed Separation Logic) [289], GPS (Ghost states, Protocols, and Separation) [286], and FSL (Fenced Separation Logic) [112; 113]. All three are based on the axiomatic C memory model [50], and are mechanised in Rocq (formerly Coq). The case-studies of these logics include simple litmus tests, while GPS in particular was applied to more realistic concurrent data structures such as the Michael-Scott queue, the circular buffer, and the bounded ticket lock.
@@
The Iris framework [163; 162] uses an underlying operational semantics to unify various concurrent separation logics from the setting of Sequential Consistency. Iris is mechanised in Rocq, which is also used by its derivatives listed below.
Iris was extended to the weak memory setting by combining with GPS. The results are iGPS [164] and iRC11 [102]; see also [275]. The case-studies of iGPS are the same as those of GPS, while the case-studies of iRC11 are drawn from a library of the Rust programming language.
Cosmo [217] is another CSL based on Iris for reasoning about multicore OCaml and its operationally defined memory model. The case-studies are a spin lock, a ticket lock, and Peterson’s algorithm. The sequel [216] then applied Cosmo to verify a concurrent bounded queue. See [121] for an alternative approach to CSL.
@@
There is a rely-guarantee logic [93] for multicopy-atomic hardware memory models, using a reordering formalism [89; 85]. The case-study is the Chase-Lev double-ended queue for Arm, with the proof mechanised in Isabelle/HOL. The same authors later extended their framework to handle memory models without multicopy atomicity [94], and verified Peterson’s algorithm in Isabelle/HOL. The CLH fine-grained lock, as used in the seL4 microkernel [169], is verified for Arm using a related rely-guarantee-based approach in [88].
@@ 8.1. Model Checkers
| CBMC extended [31] | BMC | x86, Power |
| Herd [27] | simulator | parametric (Cat language) |
| SATCheck [104] | SMC | TSO |
| CDSCheck [222; 223] | SMC | C |
| Dartagnan [142] | BMC | parametric (Cat language) |
| GenMC [175] | SMC | parametric |
| Zord [119] | BMC | TSO, PSO |
| McSPIN [14] | BMC | parametric |
| Nidhugg [9; 2] | SMC | TSO, PSO, Power |
In the weak-memory realm, most model checkers employ either Stateless Model Checking [129] or Bounded Model Checking [83]. Here, we classify the prominent model checkers into those based on axiomatic formalisms (Section 5), and, less commonly, those based on operational formalisms (Section 6.2). Table 6 summarises these model checkers, together with the memory models they support. In the text below, we also cover the case-studies to which these model checkers have been applied.
@@
Model checking under weak memory models is mostly based on axiomatic formalisms. The most prominent axiomatic formalism is the Cat language, with the associated Herd tool [36]. Herd simulates the possible behaviours of assembly programs, under memory models specified in Cat. Herd and Cat have been applied to thousands of litmus tests to check whether a specification of a memory model does agree with the hardware behaviour. They were used by Arm as the basis for developing the specifications of its memory models [27]. Herd is included in the “diy tool-suite” [39], which includes facilities for automatically generating litmus tests and running them directly on hardware.
The CBMC bounded model checker [84] was extended to handle the memory models of x86 and Power [31]. Apart from litmus tests, the case-studies include components of system software, such as a queue mechanism of the Apache HTTP server, the worker synchronisation mechanism in PostgreSQL, and the read-copy-update mechanism of the Linux kernel.
Dartagnan [235; 125; 142] is a bounded model checker, where users specify memory models in the Cat language. The benchmarks are several mutual exclusion algorithms, including Dekker’s algorithm, Peterson’s algorithm and the qspinlock of the Linux kernel [141; 229].
Another recent bounded model checker is Zord [119], which is based on a novel axiomatic formalism that supports TSO and PSO, and has been applied to the SV-COMP 2020 benchmarks.
@@
McSPIN [14] is an extension to the SPIN model checker, and users specify memory models by constraining the operational model in McSPIN. The case-studies include litmus tests from the documentation of Unified Parallel C and Itanium, as well as the double-checked locking algorithm and Dekker’s algorithm.
Nidhugg [9; 2] uses stateless model checking based on a specialised operational formalism of memory models. The memory models supported are TSO, PSO, Power, and a subset of Arm. Nidhugg accepts standard C input, and has been applied to classical benchmarks, such as Dekker’s, Lamport’s and Peterson’s mutual exclusion algorithms.
@@ 8.2. Algorithmic Problems
At the heart of model checkers are computational problems for deciding whether a program satisfies certain properties. The decidability and computational complexity of these problems are substantial topics in their own right [59]. This section surveys the works that address these problems under weak memory models, with the results summarised in Table 7.
@@
| SC | NP-com [128] | FPT [80] | |||
| Causal | NP-hard [61; 124] | P [61] | |||
| TSO | NP-hard [124] | FPT [80] | PSPACE-com [62] | dec [42] | PSPACE-com [12] |
| PSO | NP-hard [124] | FPT [80] | dec [42] | ||
| RMO | NP-hard [124] | FPT [80] | undec [42] | ||
| Power | PSPACE-com [105] | ||||
| RA | FTP [285] | undec [3] | PSPACE-com [176] |
@@
Gibbons and Korach [128] first showed that consistency checking under Sequential Consistency is NP-complete. Later, Cantin et al. [72; 73] showed that it is NP-complete to check if a concurrent program yields a coherent execution (Definition 5.10). As a corollary, consistency checking is NP-hard under most memory models, since most memory models satisfy the coherence property. Indeed, Furbach et al. [124] established the NP-hardness of consistency checking under the memory models covered by Steinke and Nutt [272].
@@
Under RMO the reachability problem is undecidable, but under TSO and PSO it is decidable, albeit beyond the primitive recursive (PR) class and has a very high computational cost [42]. In light of this computational hardness, there were works that attempted to trade completeness with efficiency under TSO [60].

### v2

@@ 6.3. Axiomatic Formalisms
One of the most successful approaches to formalising memory models is via constraints on the allowed behaviours. This is referred to as the axiomatic (or declarative) approach in the literature [Alglave et al., 2014b; Sindhu et al., 1992]. Given a set of memory events (where each read has a return value), one starts with all possible permutations, and retains only those permutations that satisfy the axioms of the particular memory model (conceptually the opposite of first considering the set of behaviours allowed under Sequential Consistency and then expanding that set to capture the weak behaviours).
Due to its preciseness and relative conciseness, the axiomatic approach is widely adopted for the major memory models, e.g., x86-TSO [Owens et al., 2009], Arm [Alglave et al., 2021], RISC-V [RISC-V Foundation, 2019], Java [Manson et al., 2005], and C [Batty et al., 2011].
Some of the most influential tools in this area are the domain-specific language Cat and its associated litmus-test simulator Herd, introduced in [Alglave et al., 2014b]. Cat and Herd are utilised by Arm to formalise its memory model [Alglave et al., 2021], and Cat is also used in a number of model checkers for expressing other memory models (see Section 8.1). An axiomatic formalisation is leveraged by the Check suite of tools and techniques for investigating correctness and security of microarchitectural implementations [Lustig et al., 2014; Manerkar et al., 2018; Trippel et al., 2019; Manerkar et al., 2015; Lustig et al., 2016; Trippel et al., 2018].
The downside of axiomatic formalisms is that they are based on global reasoning about completed executions, where every has a value, and the relations and are supplied (see Section 5). As such, axiomatic formalisms tend not to be suited for reasoning about programs compositionally; that is, one cannot form the set of behaviours of a system by considering individual processes in isolation. The purely axiomatic approach has less applicability to theorem proving (however, see [Alglave and Cousot, 2017; Hammond et al., 2024]), so operational semantics (see Section 6.2) are sometimes developed in tandem, often with an underlying axiomatic flavour [Doherty et al., 2022; Kang et al., 2017; Pulte et al., 2019; Crary and Sullivan, 2015].
@@ 6.4. Event Structures and Pomsets
Both event structures [Winskel, 1987] and pomsets [Pratt, 1986] were invented in the 1980s as mathematical models of general concurrency. An early use of event structures for weak memory models is by Cenciarelli et al. [2007] in 2007, who described a small fragment of the Java memory model by combining event structures with operational and axiomatic techniques. Then from 2016 onwards, there have been many attempts to define memory models using related ideas [Castellan, 2016; Jeffrey and Riely, 2016; Pichon-Pharabod and Sewell, 2016; Chakraborty and Vafeiadis, 2019; Paviotti et al., 2020; Jagadeesan et al., 2020; Jeffrey et al., 2022; Moiseenko et al., 2022]. The Weakestmo Semantics and its extension [Chakraborty and Vafeiadis, 2019; Moiseenko et al., 2022] are multi-execution, i.e., they encode multiple execution graphs in a single event structure (cf. the Promising Semantics Section 6.2).
@@ 6.5. Program Transformations and Instruction Reorderings
A number of early works [Adve and Gharachorloo, 1996; McKenney, 2005a, b] specified weak memory models in terms of which pairs of instructions can be reordered. This aligns with the workings of the processor pipeline (Definition 2.1), and is often the level at which processors are described informally. The first formal memory-model framework based on instruction reordering was developed by Arvind and Maessen [2006], who included store-atomicity as an extra criterion.
Possible instruction reordering is important in the context of compiler-related aspects of weak memory models. Common program transformations (including instruction reordering) performed by optimising compilers are not always correct under weak memory models [Ševčík and Aspinall, 2008; Vafeiadis et al., 2015]. To formulate what it means for a program transformation to be “correct” under a memory model, one typically develops some trace-based denotational semantics, and a transformation is then correct if the denotation of the transformed program is (a subset of) the denotation of the original program. Works in this line of research include Ševčík and Aspinall [2008] and Burckhardt et al. [2010] for the Java memory model, and Dodds et al. [2018] and Dvir et al. [2024] for the C memory model and its RA fragment.
Instruction reordering was also the basis in [Colvin, 2021], where the memory model is encoded as a parameter to sequential composition on the level of the program text. Weak memory effects are explained in terms of the standard operators for (no-parameter) sequential composition and parallel composition, via a set of transformation rules on the instructions, in the spirit of the Concurrent Kleene Algebra [Hoare et al., 2009; Hayes et al., 2016]. For example, instantiating the SB litmus test on an x86 (TSO) machine, one can derive the following.
@@
This reasoning framework has been used as the basis for addressing security vulnerabilities related to speculative execution [Colvin and Winter, 2020] and information flow [Smith et al., 2019]. By transforming a program under a weak memory model to a program under Sequential Consistency, standard reasoning methods and tools become amenable, saving the need for specialised logics and techniques. However, in contrast to timestamped messages, these program-level transformations cannot explain non-multicopy-atomic behaviours [Lahav and Vafeiadis, 2016].
@@ 7. Program Logics and mechanised support
There are three prominent families of program logics for concurrent programs that have been adapted to reason about weak memory effects: Owicki-Gries (OG) [Owicki and Gries, 1976], Rely-Guarantee (RG) [Jones, 1983a, b], and Concurrent Separation Logic (CSL) [Ishtiaq and O’Hearn, 2001; Reynolds, 2002; Brookes and O’Hearn, 2016]. This section surveys these adaptations; a summary is given in Table 5.
@@
| OG/RG | Axiomatic | C fragment | RCU | [Lahav and Vafeiadis, 2015] |
| RG | Operational | x86-TSO | Simpson’s | [Ridge, 2010] |
| RG | Operational | (parameterised) | litmus tests | [Abe and Maeda, 2016] |
| OG | View-based | C fragment | Peterson’s, RCU | [Dalvandi et al., 2020, 2022; Doherty et al., 2022; Wright et al., 2021] |
| RG | potential-based | Causal Consistency | litmus tests, Peterson’s | [Lahav et al., 2023] |
| CSL | Axiomatic | C fragment | MS queue and others | [Doko and Vafeiadis, 2016, 2017; Vafeiadis and Narayan, 2013; Turon et al., 2014] |
| CSL | Operational | C fragment | MS queue and others | [Kaiser et al., 2017] |
| CSL | Operational | Rust | Rust library | [Dang et al., 2019] |
| CSL | Operational | OCaml | locks, bounded queue | [Mével et al., 2020; Mével and Jourdan, 2021] |
| RG | Reordering | hardware models | Chase-Lev, Peterson’s | [Coughlin et al., 2021, 2022] |
| Invariant | Axiomatic | (parameterised) | PostgreSQL | [Alglave et al., 2013b; Alglave and Cousot, 2017] |
| Invariant | Operational | C fragment | Peterson’s | [Doherty et al., 2019] |
@@
Based on an axiomatic formulation of the release-acquire fragment of the C memory model, a rely-guarantee logic was developed and applied to the read-copy-update synchronisation mechanism [Lahav and Vafeiadis, 2015].
@@
A rely-guarantee logic on the x86 assembly instructions was developed in [Ridge, 2010], based on the operational formulation of x86-TSO [Owens et al., 2009]. This logic was applied to Simpson’s four-slot algorithm, and mechanised in HOL. The rely-guarantee logic of [Abe and Maeda, 2016] works on general memory models that are specified operationally, with case-studies performed on litmus tests.
Another line of research [Dalvandi et al., 2020, 2022] devised a view-based formalisation (see Section 6.2) of the release-acquire-relaxed fragment of the C memory model, and built an Owicki-Gries logic based on this formulation. The main case-studies, carried out in Isabelle/HOL, are Peterson’s algorithm and the read-copy-update mechanism; related works in this line of research include [Doherty et al., 2022; Wright et al., 2021].
Similar to view-based semantics is the “potential-based” semantics, which has been used as the basis of a rely-guarantee logic for Causal Consistency [Lahav et al., 2023].
@@
One line of research consists of the following three generations of logics: RSL (Relaxed Separation Logic) [Vafeiadis and Narayan, 2013], GPS (Ghost states, Protocols, and Separation) [Turon et al., 2014], and FSL (Fenced Separation Logic) [Doko and Vafeiadis, 2016, 2017]. All three are based on the axiomatic C memory model [Batty et al., 2011], and are mechanised in Rocq (formerly Coq). The case-studies of these logics include simple litmus tests, while GPS in particular was applied to more realistic concurrent data structures such as the Michael-Scott queue, the circular buffer, and the bounded ticket lock.
@@
The Iris framework [Jung et al., 2015, 2018] uses an underlying operational semantics to unify various concurrent separation logics from the setting of Sequential Consistency. Iris is mechanised in Rocq, which is also used by its derivatives listed below.
Iris was extended to the weak memory setting by combining with GPS. The results are iGPS [Kaiser et al., 2017] and iRC11 [Dang et al., 2019]; see also [Svendsen et al., 2018]. The case-studies of iGPS are the same as those of GPS, while the case-studies of iRC11 are drawn from a library of the Rust programming language.
Cosmo [Mével et al., 2020] is another CSL based on Iris for reasoning about multicore OCaml and its operationally defined memory model. The case-studies are a spin lock, a ticket lock, and Peterson’s algorithm. The sequel [Mével and Jourdan, 2021] then applied Cosmo to verify a concurrent bounded queue. See [Ferreira et al., 2010] for an alternative approach to CSL.
@@
There is a rely-guarantee logic [Coughlin et al., 2021] for multicopy-atomic hardware memory models, using a reordering formalism [Colvin and Smith, 2018; Colvin, 2021]. The case-study is the Chase-Lev double-ended queue for Arm, with the proof mechanised in Isabelle/HOL. The same authors later extended their framework to handle memory models without multicopy atomicity [Coughlin et al., 2022], and verified Peterson’s algorithm in Isabelle/HOL. The CLH fine-grained lock, as used in the seL4 microkernel [Klein et al., 2009], is verified for Arm using a related rely-guarantee-based approach in [Colvin et al., 2024].
@@
There are other proof methods for invariant-based reasoning [Alglave et al., 2013b; Alglave and Cousot, 2017; Doherty et al., 2019]. The case studies in [Alglave et al., 2013b; Alglave and Cousot, 2017] are drawn from PostgreSQL, whereas in [Doherty et al., 2019] the case study is Peterson’s mutual exclusion algorithm.
@@ 8.1. Model Checkers
| CBMC extended [Alglave et al., 2013a] | BMC | x86, Power |
| Herd [Alglave et al., 2021] | simulator | parametric (Cat language) |
| SATCheck [Demsky and Lam, 2015] | SMC | TSO |
| CDSCheck [Norris and Demsky, 2013, 2016] | SMC | C |
| Dartagnan [Haas et al., 2022] | BMC | parametric (Cat language) |
| GenMC [Kokologiannakis and Vafeiadis, 2021] | SMC | parametric |
| Zord [Fan et al., 2023] | BMC | TSO, PSO |
| McSPIN [Abe and Maeda, 2017] | BMC | parametric |
| Nidhugg [Abdulla et al., 2016, 2017] | SMC | TSO, PSO, Power |
In the weak-memory realm, most model checkers employ either Stateless Model Checking [Godefroid, 1997] or Bounded Model Checking [Clarke et al., 2001]. Here, we classify the prominent model checkers into those based on axiomatic formalisms (Section 5), and, less commonly, those based on operational formalisms (Section 6.2). Table 6 summarises these model checkers, together with the memory models they support. In the text below, we also cover the case-studies to which these model checkers have been applied.
@@
Model checking under weak memory models is mostly based on axiomatic formalisms. The most prominent axiomatic formalism is the Cat language, with the associated Herd tool [Alglave et al., 2014b]. Herd simulates the possible behaviours of assembly programs, under memory models specified in Cat. Herd and Cat have been applied to thousands of litmus tests to check whether a specification of a memory model does agree with the hardware behaviour. They were used by Arm as the basis for developing the specifications of its memory models [Alglave et al., 2021]. Herd is included in the “diy tool-suite” [Arm Ltd., 2024], which includes facilities for automatically generating litmus tests and running them directly on hardware.
The CBMC bounded model checker [Clarke et al., 2004] was extended to handle the memory models of x86 and Power [Alglave et al., 2013a]. Apart from litmus tests, the case-studies include components of system software, such as a queue mechanism of the Apache HTTP server, the worker synchronisation mechanism in PostgreSQL, and the read-copy-update mechanism of the Linux kernel.
Dartagnan [Ponce de León et al., 2018; Gavrilenko et al., 2019; Haas et al., 2022] is a bounded model checker, where users specify memory models in the Cat language. The benchmarks are several mutual exclusion algorithms, including Dekker’s algorithm, Peterson’s algorithm and the qspinlock of the Linux kernel [Haas et al., 2023; Paolillo et al., 2021].
Another recent bounded model checker is Zord [Fan et al., 2023], which is based on a novel axiomatic formalism that supports TSO and PSO, and has been applied to the SV-COMP 2020 benchmarks.
Meanwhile, there are a few stateless model checkers that are based on axiomatic formalisms. For instance, CDSCheck [Norris and Demsky, 2013, 2016] is based on the axiomatic C memory model [Batty et al., 2011], and uncovered an error in an implementation of the Chase-Lev double-ended queue.
There is also the line of stateless model checkers HMC [Kokologiannakis and Vafeiadis, 2020], GenMC [Kokologiannakis and Vafeiadis, 2021] and WMC [Moiseenko et al., 2022]. They have been applied to synthetic litmus tests, mutual exclusion algorithms, and concurrent data structures (Chase-Lev double-ended queue, Treiber stack, and Michael-Scott queue).
@@
McSPIN [Abe and Maeda, 2017] is an extension to the SPIN model checker, and users specify memory models by constraining the operational model in McSPIN. The case-studies include litmus tests from the documentation of Unified Parallel C and Itanium, as well as the double-checked locking algorithm and Dekker’s algorithm.
Nidhugg [Abdulla et al., 2016, 2017] uses stateless model checking based on a specialised operational formalism of memory models. The memory models supported are TSO, PSO, Power, and a subset of Arm. Nidhugg accepts standard C input, and has been applied to classical benchmarks, such as Dekker’s, Lamport’s and Peterson’s mutual exclusion algorithms.
@@ 8.2. Algorithmic Problems
At the heart of model checkers are computational problems for deciding whether a program satisfies certain properties. The decidability and computational complexity of these problems are substantial topics in their own right [Bouajjani, 2024]. This section surveys the works that address these problems under weak memory models, with the results summarised in Table 7.
@@
| SC | NP-com [Gibbons and Korach, 1997] | FPT [Chini and Saivasan, 2020] | |||
| Causal | NP-hard [Bouajjani et al., 2017; Furbach et al., 2015] | P [Bouajjani et al., 2017] | |||
| TSO | NP-hard [Furbach et al., 2015] | FPT [Chini and Saivasan, 2020] | PSPACE-com [Bouajjani et al., 2011] | dec [Atig et al., 2010] | PSPACE-com [Abdulla et al., 2020] |
| PSO | NP-hard [Furbach et al., 2015] | FPT [Chini and Saivasan, 2020] | dec [Atig et al., 2010] | ||
| RMO | NP-hard [Furbach et al., 2015] | FPT [Chini and Saivasan, 2020] | undec [Atig et al., 2010] | ||
| Power | PSPACE-com [Derevenetc and Meyer, 2014] | ||||
| RA | FTP [Tunç et al., 2023] | undec [Abdulla et al., 2019] | PSPACE-com [Krishna et al., 2022] |
@@
Gibbons and Korach [1997] first showed that consistency checking under Sequential Consistency is NP-complete. Later, Cantin et al. [Cantin et al., 2003, 2005] showed that it is NP-complete to check if a concurrent program yields a coherent execution (Definition 5.10). As a corollary, consistency checking is NP-hard under most memory models, since most memory models satisfy the coherence property. Indeed, Furbach et al. [2015] established the NP-hardness of consistency checking under the memory models covered by Steinke and Nutt [2004].
@@
Under RMO the reachability problem is undecidable, but under TSO and PSO it is decidable, albeit beyond the primitive recursive (PR) class and has a very high computational cost [Atig et al., 2010]. In light of this computational hardness, there were works that attempted to trade completeness with efficiency under TSO [Bouajjani et al., 2015].